Solve the equation.
15r^2 = 3r

Answer by tutorcecilia(2152)

15r^2 = 3r
(3r)(5r-1)=0
3r=0
3r/3=0/3
r=0
and
5r-1=0
5r=1
5r/5=1/5
r=1/5
The solutions are r=0
and r=1/5